OpenStudy (anonymous):

3x+(1-x)4=2(4-x) I got 'No answer' but I'm not sure if it's correct =/

HanAkoSolo (jamierox4ev3r):

3x+4-4x=8-2x -x+4=8-2x -4 +2x -x=8 x=-8

OpenStudy (anonymous):

I'm kinda confused about the (1-x)4 thing....

HanAkoSolo (jamierox4ev3r):

its just a different way of writing a problem of distribution. Just think about the four being on the left side (the opposite side)

OpenStudy (johnweldon1993):

Right...but most times on this site...people write it like that when they mean it to be an exponent... just want to make sure...

HanAkoSolo (jamierox4ev3r):

ah i see

HanAkoSolo (jamierox4ev3r):

@KawasumiKimiko

OpenStudy (anonymous):

yeah because i thought that (1-x)4 was equal to 4(1-x), which it is

HanAkoSolo (jamierox4ev3r):

unless, as johnweldon pointed out, it could be an exponent, but idk...

OpenStudy (anonymous):

which is why I was a little hesitant of writing the answer to the question

OpenStudy (johnweldon1993):

@KawasumiKimiko is the equation: \[3x+(1-x)^4=2(4-x)\] or \[3x+(1-x)4=2(4-x)\] ?

HanAkoSolo (jamierox4ev3r):

@KawasumiKimiko is the 4 at the end of the parentheses an exponent or is it just a term that can be distributed into the parentheses?

OpenStudy (anonymous):

@johnweldon1993 the second one.

OpenStudy (johnweldon1993):

Ahh okay...the @Jamierox4ev3r has done it *almost correctly* above 3x+(1-x)4=2(4-x) distribute the 4 into the parenthesis...and the 2 into the other parenthesis 3x+4-4x=8-2x subtract 4 from both sides 3x+4-4x=8-2x -4 -4 3x -4x=4 -2x simplify the left hand side 3x - 4x = -1x or just -x...so -x = 4 - 2x Add 2x to both sides -x = 4 - 2x +2x +2x x = 4 Just remember @Jamierox4ev3r ....what you do to 1 side...you do to the other...so when you subtract 4...you subtract it from BOTH sides...and when you add 2x...you add 2x to BOTH sides...

OpenStudy (anonymous):

Thank you guys so so so so much! ^^ sorry for all the trouble :3

OpenStudy (johnweldon1993):

No problem! :)